OpenStudy (anonymous):

Find the number of complex roots of x3 + 4x2 + 3x = 0. What are the roots of the function?

OpenStudy (anonymous):

first you have to take out one x from the equation like this x(x^2+4x+3)=0 then solve the equation x(x^2+4x+3)=0 so we have x(x+1)(x+3)=0, then we have the roots of this function like this x=0, x= -1, x=-3